# The Open Journal of Astrophysics

### v1

@@ Abstract
Super-resolution (SR) models in cosmological simulations use deep learning (DL) to rapidly supplement low-resolution (LR) runs with statistically correct, fine details. The SR technique preserves large-scale structures by conditioning on a low-resolution (LR) version of the simulation. On smaller scales, the generative deep learning (DL) process is stochastic, resulting in numerous possible SR realizations, each with unique small-scale structures. Validation of SR then relies on making sure that a specific statistic of interest is accurately reproduced by comparing SR and high resolution (HR) runs. In this study, we develop an emulator designed to reproduce the individual small-scale structures of an HR simulation as closely as possible. We process an SR realization alongside a specific High-Resolution Initial Condition (HRIC), transforming the SR output to emulate the results of a full simulation with that HRIC. By comparing visualizations, individual halo measures and cross-correlating Fourier modes we show that the emulated SR runs closely align with the corresponding HR simulation, even on length scales an order of magnitude smaller than the LR run. Additionally, small halos are trained to match the HR simulation, and the subhalo mass function is more accurately reproduced. These results show the promise of this method for generating numerous fast and accurate simulations and mock observations for large galaxy surveys.
@@ 1 Introduction
Cosmological N-body simulations have served as powerful computational tools since the 1960s (von Hoerner, 1960; Aarseth, 1963) for solving the non-linear dynamics intrinsic to cosmic structure formation. High-resolution (HR) simulations employing state-of-the-art, massively-parallel codes, such as MP-Gadget, evolve a large number of particles under gravity, tracking the formation of structures on the order of galaxy size and larger. These simulations demand significant time and computational resources, forcing researchers to make trade-offs between resolution and size of simulation volumes. Hydrodynamic simulations, such as Illustris-TNG (Springel et al., 2018), SIMBA (Davé et al., 2019) and ASTRID (Bird et al., 2022; Ni et al., 2022), which model the baryonic matter as a separate component, require even more computational investment compared to gravity-only N-body simulations. Consequently, researchers are often compelled to reduce resolution due to the intensive computational demands or, alternatively, select smaller volumes to allow for finer scales. For example, the small volume high-resolution FIRE suite (Hopkins et al., 2014) and the ABACUS summit (Maksimova et al., 2021) runs optimized for large-scale structure. In the present paper we explore the use of Machine Learning (ML) techniques (Yue et al., 2016) to sidestep these limitations by vastly accelerating simulations by many orders of magnitude. We anticipate that these hybrid N-body-ML simulations will make certain tasks much easier (e.g., making mock catalogs for large galaxy surveys).
@@
In recent years, ML has become a promising tool for physicists, particularly the use of Deep Learning (DL) (Dvorkin et al., 2022). ML models have been developed to predict different baryonic properties from dark matter-only simulations, including the galaxy distribution (Modi et al., 2018; Zhang et al., 2019), the tSZ effect (Tröster et al., 2019), the 21 cm emission distribution (Wadekar et al., 2020), and stellar maps (Dai & Seljak, 2021). Predictions of nonlinear structure from linear cosmological initial conditions  (He et al., 2019; Berger & Stein, 2019; Bernardini et al., 2020; Alves de Oliveira et al., 2020) and from the dark matter density field (Kodi Ramanah et al., 2020a) have proved that DL based Neural Networks (NNs) can capture highly nonlinear structures. Training large NNs requires extensive, high-quality datasets. CAMELS (Cosmology and Astrophysics with MachinE Learning Simulations, Villaescusa-Navarro et al., 2020), a large-scale project with a training set of over 4000 hydrodynamical simulations run with different hydrodynamic solvers and subgrid models for galaxy formation. A prime purpose of this set is to investigate the interplay of baryonic effects and cosmology, using machine learning techniques.
@@
In a quest to achieve faster results comparable in accuracy to full HR N-body simulations, researchers have applied "Super-Resolution" (SR) techniques to lower resolution cosmological simulations (see e.g., Li et al. 2021 and other papers in this series). SR in general describes the situation where information is added below the resolution scale of initially Low Resolution (LR) data such as a photograph (e.g., in Johnson et al. 2016) to create an HR version. However, SR remains a challenging and ill-posed problem in ML, because an infinite number of HR results can be downsampled to the same LR input, meaning the SR model learns a conditional probabilistic distribution of HR simulations based on a single LR simulation. This ‘one-to-many’ regime can be useful when many simulations or mock observations are needed for accurate and precise cosmological parameter inference. It can however make validation difficult, as it can only be done in a statistical sense by averaging over ensembles.
@@
Generative Adversarial Network (GAN) (Goodfellow et al., 2014) architectures have already proven to be effective in SR tasks (Wang et al., 2019) and have been applied to the generation of 2D images of cosmic webs (Rodríguez et al., 2018), 3D matter density fields (Perraudin et al., 2019) and cosmological mass maps (Perraudin et al., 2020). GANs consist of two separate NNs: a generator, which generates candidates, and a discriminator, which provides feedback by comparing real data with generated candidates. In previous work, an SR framework which applies a GAN approach directly to $N$-body simulation particles has been explored, covering from displacement only (Li et al., 2021) to full 6-D phase space (Ni et al., 2021), and which can also conditioned on redshift (Zhang et al., 2023) (hereafter Paper I, Paper II and Paper III). The output of this framework can be treated and analyzed in the same manner as full $N$-body simulation data. The same framework has also been used for a fuzzy dark matter cosmology (Sipp et al., 2023). In Paper I, II and III, we trained our GAN based framework to perform a stochastic mapping from LR input to SR output, where stochasticity is injected during the generation process. These models aim to learn a probability distribution of HR outputs conditioned on LR inputs. By sampling this distribution using a special technique, as described in Zhang et al. (2023), we can carefully control the randomness in the SR field, allowing us to generate simulations with outputs at any redshift desired. Instead of checking that summary statistics of the SR model are within expected bounds, an SR run that reproduces the individual structures of the HR run would (at least down to some length or mass scale) be a valid cosmological simulation, one that could be used to compute any summary statistic. This is one of the goals of this paper.
Recently, denoising diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have become state-of-the-art, outperforming GANs for some image generation tasks and also being used in SR. It has been shown that super-resolving a 3D density field conditioned on an LR density field (Rouhiainen et al., 2023) using a diffusion model has comparable performance to GAN results and does not suffer from unstable training and mode collapse. Several approaches employing diffusion models demonstrate the potential of this method, such as in 2D galaxy images (Smith et al., 2022) and in 2D simulations (Schanz et al., 2023) using filter-boosted tricks. However, their work is limited to 2D cases only.
@@
In this work, we introduce an emulator framework which can adjust the SR outputs to their HR counterparts by conditioning on specific high-resolution initial conditions (HRICs). These specific HRICs have the same large-scale modes as the LR $N$-body simulations that are used to generate SR outputs. Other studies have employed emulators, such as Kodi Ramanah et al. (2020b), which combines LR and HRICs for super-resolution tasks on 3D density fields. Notably, no stochasticity is injected into the model, similar to our approach. Additionally, Jamieson et al. (2023) uses a neural network to map the linear approximation at $z=0$ to real $N$-body simulations with various cosmological parameters, using a Lagrangian description. To achieve accurate SR emulation of the HR simulations down to small-scales, we find that using adversarial training techniques are crucial for correcting small-scale discrepancies. We explore emulators trained with this technique (hereafter referred to as "Emulator"), and compare them with our previous work in Zhang et al. (2023) (hereafter referred to as "SR").
@@
The paper is structured as follows. In Section 2, we discuss the generation of data sets and summarize the details of our neural networks and the training process. In Section 3 we show the performance of our emulator, specifically visual validation in 3.1, Fourier mode cross-correlations in 3.2 and halo catalog in 3.3. In Section 4, we discuss our findings, and we summarize our results and conclude in Section 5. We have also included the results of training without discriminator guidance in Appendix A.
@@ 2.1 Dataset
The $N$-body simulation is a widely used tool for understanding the nonlinear evolution of cosmic structures. Given a fixed set of cosmological parameters (in this work, we use the WMAP9 cosmology, as prescribed by Hinshaw et al., 2013), this method involves dividing the mass distribution into $N$ equal-mass particles. We run both LR and HR simulations using MP-Gadget [^1] https://github.com/MP-Gadget/MP-Gadget, a highly parallel code that solves the gravitational force using the TreePM method (Bagla, 2002). Long-range forces are calculated on a particle mesh in Fourier space, and short-range forces are computed through a hiearchical tree structure. Our LR and HR simulation tasks follow the Lagrangian description, with particles are tracers of both displacement and velocity fields. The displacement of an individual particle is:
@@
$\mathbf{v}_{i}\equiv\dot{\mathbf{x}_{i}}=\dot{\mathbf{d}_{i}},$ (2)
where the dots denote time derivatives. The initial conditions for our simulations are generated at $z=99$ using a power spectrum calculated by CLASS (Lesgourgues, 2011). First-order Lagrangian perturbation theory (Zel’dovich, 1970; Crocce et al., 2006) is used to initialize particle positions and velocities. The definition of initial displacement $\mathbf{d}_{init,i}$ and velocity $\mathbf{v}_{init,i}$ for particles is similar to Eq. 1 and 2. Cosmological parameters follow the WMAP9 cosmology (Hinshaw et al., 2013), with matter density $\Omega_{\rm m}=0.2814$, dark energy density $\Omega_{\Lambda}=0.7186$, baryon density $\Omega{\rm b}=0.0464$, power spectrum normalization $\sigma_{8}=0.82$, spectral index $n_{s}=0.971$, and Hubble parameter $h=0.697$. The mean spatial separation of the dark matter particles is used to determine the gravitational softening length, which is set to $1/30$ of this value. Our training dataset is comprised of 8 dark matter only simulations, each with LR and HR simulation pairs. Each LR-HR simulation pair has $64^{3}$ and $512^{3}$ particles, respectively, with box size $100h^{-1}{\rm Mpc}$. The mass resolution is $m_{\mathrm{DM}}=2.98\times 10^{11}{h^{-1}M_{\odot}}$ for LR, and $m_{\mathrm{DM}}=5.8\times 10^{8}{h^{-1}M_{\odot}}$ for HR. Each pair has a unique random seed for its initial condition, so the initial conditions differ for eight simulations. In addition, we also create a test dataset with 3 LR and HR simulation pairs. Each pair consists of 60 snapshots from $z=1$ to $z=0$, with the same cosmological parameters as the training set but different initial condition random seeds. During data pre-processing, we first divide the particle displacement and velocity fields into cubical chunks. We then concatenate the displacement field and velocity field channel-wise since they share the same spatial resolution. This process is applied similarly to the HRICs.
@@ 2.2 Model Architecture and General Training Process
Figure 1: Upper Panel : Schematic plot of the main components of the model and the entire training process. Lower Panel : Architecture of our U-Net model and detailed structure of our Residual Block. The Res up Block or Res Down Block blocks correspond to the up-sampling or down-sampling blocks in the detailed structure, while the Res Blocks do not have up/down sampling blocks.
@@
U-Net is a universal architecture that was first designed for semantic segmentation (Ronneberger et al., 2015). It consists of an encoder and a decoder, with a bottleneck between them. The encoder progressively downsamples the input towards the bottleneck, while the decoder upsamples the data from bottleneck back to the original resoltion. "Skip-connections" link the encoder and decoder at corresponding resolutions. U-Nets have also been applied to super-resolution/style transfer tasks (Johnson et al., 2016), image to image translation, image uncropping and more (see discussion in Saharia et al., 2021, for a review). Recent diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) also use U-Nets for denoising.
@@
The lower panel of Figure 1 illustrates the detailed architecture of our U-Net model. The encoder part extracts features from the input and downsamples them towards the bottleneck. This forces the model to learn the key information from the input and embed it into a low-dimensional space. Starting from the input resolution, the decoder downsamples the learned feature until it reaches the bottleneck blocks. Each decoder block at the same resolution consists of two residual blocks. We follow BigGAN’s (Brock et al., 2018) residual block structure since it has been proven to have good performance and can be scaled to different depths. All residual blocks contain two branches, one “skip” branch with a 3D size 1 convolutional block, the other “main” branch, which first normalizes the input with Group Normalization (Wu & He, 2018), then passes through a Sigmoid Linear Unit (SiLU) activation function (Elfwing et al., 2017). The second group normalization layer is applied to normalize the output from a 3D size 3 convolutional block. Finally, after one more activation function and one 3D size 3 convolutional block, the output from the “main” branch is added to the output from the “skip” branch. The Up/Down layer denotes the upsampling or downsampling procedure, only applied in Res Up/Down Blocks. The decoder part decodes and upsamples the features from the bottleneck and the “skip” connections from the encoder. The information from the encoder is passed to the decoder at multiple resolution levels, to maintain information at different levels and avoid vanishing gradient issues. No stochasticity is injected in this model.
@@
The goal of our model is to adjust the fields output by a stochastic SR model to those that corresponds as exactly as possible to the output of a specific HR simulation by conditioning on the HRICs. Due to the limitations of GPU memory, we first split the SR field and HRICs into equal size chunks, then concatenate them in a channel-wise fashion since they have the same spatial dimensions. Both chunks pass through the encoder, bottleneck and decoder steps together, and are then compared to the real HR simulations generated from the same HRICs. To evaluate our emulator’s output, we use the Charbonnier loss function (Barron, 2017), which is defined as
@@
Using feedback from discriminator (i.e. adversarial loss) is common in image generation and image to image translation tasks. It is used to enhance perceptual quality of the output. We use the same definition of adversarial loss $\mathcal{L}_{Adv}$ as in previous SR works. The total loss for our emulator network is:
@@
The parameter $\lambda$ is a weight to balance the contribution of the different loss terms, and the value of $\lambda$ is set empirically to $3\times 10^{-2}$. We have settled on this value after experimentation to balance the accuracy and perceptual quality of the outputs.

### v2

@@ Abstract
Super-resolution (SR) models in cosmological simulations use deep learning (DL) to rapidly enhance low-resolution (LR) runs with statistically correct fine details. These models preserves large-scale structures by conditioning on an LR version of the simulation. On smaller scales, the generative process is inherently stochastic, producing multiple possible SR realizations with distinct small-scale structures. Validation of reconstructed SR runs from LR simulations requires ensuring that specific statistics of interest are accurately reproduced by comparing SR outputs with target high resolution (HR) runs. In this study, we develop an emulator designed to reproduce the small-scale structures of target HR simulation with high fidelity. By processing an SR realization alongside the high-resolution initial condition (HRIC), we transform the SR output to emulate the result of a full simulation with that HRIC. By comparing various metrics, from visualization to individual halo measurements, we demonstrate that the emulated SR runs closely align with the target HR simulation, even at length scales an order of magnitude smaller than the corresponding LR run. These results show the potential of this method for efficiently generating accurate simulations and mock observations for large galaxy surveys.
@@ 1 Introduction
Cosmological N-body simulations have served as powerful computational tools since the 1960s (von Hoerner, 1960; Aarseth, 1963) for solving the nonlinear dynamics intrinsic to cosmic structure formation. High-resolution (HR) simulations employing state-of-the-art, massively parallel codes, such as MP-Gadget, evolve a large number of particles under gravity, tracking the formation of structures on the order of galaxy size. Hydrodynamic simulations, such as Illustris-TNG (Springel et al., 2018), SIMBA (Davé et al., 2019) and ASTRID (Bird et al., 2022; Ni et al., 2022), which model the baryonic matter as a separate component, require even more computational investment compared to gravity-only N-body simulations. These simulations require substantial time and computational resources, forcing researchers to balance resolution and simulation volume. As a result, they must often choose between reducing resolution to meet computational constraints or limiting the simulation volume to achieve finer scales. For example, the small volume, high-resolution FIRE suite (Hopkins et al., 2014) while ABACUS summit (Maksimova et al., 2021) runs are optimized for large-scale structure.
@@
In recent years, Machine Learning (ML), particularly Deep Learning (DL), has emerged as a powerful tool for physicists (Dvorkin et al., 2022). ML models have been employed to predict various baryonic properties from dark matter-only simulations, including the galaxy distribution (Modi et al., 2018; Zhang et al., 2019), the tSZ effect (Tröster et al., 2019), the 21 cm emission distribution (Wadekar et al., 2020), and stellar maps (Dai & Seljak, 2021). Predictions of nonlinear structure from linear cosmological initial conditions  (He et al., 2019; Berger & Stein, 2019; Bernardini et al., 2020; Alves de Oliveira et al., 2020) and from the dark matter density field (Kodi Ramanah et al., 2020a) have proved that DL based Neural Networks (NNs) can capture highly nonlinear structures.
Training large NNs requires extensive, high-quality datasets. An example of such a simulated dataset is CAMELS (Cosmology and Astrophysics with MachinE Learning Simulations, Villaescusa-Navarro et al., 2020), a large-scale simulation project with a training set of over 4000 hydrodynamical simulations run with different hydrodynamic solvers and subgrid models for galaxy formation. The large size and comprehensive range of parameter variations in this dataset enables investigations using machine learning techniques into the interplay of baryonic effects and cosmology.
@@
To reduce computational demand and also achieve result comparable in accuracy to full HR N-body simulations, researchers have developed Super-Resolution (SR) techniques to lower resolution cosmological simulations (see e.g., Li et al. 2021 and other papers in this series). SR in general describes the situation where information is added below the resolution scale of initially lower resolution data such as a picture (e.g., in Johnson et al. 2016) to create an HR version. However, SR remains a challenging and ill-posed problem, because an infinite number of HR results can be downsampled to the same LR input, meaning the SR model learns a conditional probabilistic distribution of HR simulations based on a single LR simulation. This ‘one-to-many’ regime can be useful when many simulations or mock observations are needed for accurate and precise cosmological parameter inference.
@@
Generative Adversarial Network (GAN) (Goodfellow et al., 2014) architectures have already proven to be effective in SR tasks (Wang et al., 2019) and have been applied to the generation of 2D images of cosmic webs (Rodríguez et al., 2018), 3D matter density fields (Perraudin et al., 2019) and cosmological mass maps (Perraudin et al., 2020). GANs consist of two separate NNs: a generator, which generates candidates, and a discriminator, which provides feedback by comparing real data with generated candidates. In previous work, an SR framework which applies a GAN approach directly to $N$-body simulation particles has been explored, covering from displacement only (Li et al., 2021) to full 6-D phase space (Ni et al., 2021), and which can also be conditioned on redshift (Zhang et al., 2023) (hereafter Paper I, Paper II and Paper III). The output of this framework can be treated and analyzed in the same manner as full $N$-body simulation data. The same framework has also been used for fuzzy dark matter cosmology (Sipp et al., 2023). In Paper I, II and III, we trained our GAN based framework to perform a stochastic mapping from LR input to SR output, where stochasticity is injected during the generation process. These models aim to learn a probability distribution of HR outputs conditioned on LR inputs. By sampling this distribution using a special technique, as described in Zhang et al. (2023), we can carefully control the randomness in the SR field, allowing us to generate simulations with outputs at any desired redshift. Instead of checking that summary statistics of the SR model are within expected bounds, an SR run that reproduces the individual structures of the HR run would (at least down to some length or mass scale) be a valid cosmological simulation, one that could be used to compute any summary statistic. This is one of the goals of this paper. Recently, denoising diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have become state-of-the-art, outperforming GANs for some image generation tasks and also being used in SR. It has been shown that super-resolving a 3D density field conditioned on an LR density field (Rouhiainen et al., 2023) using a diffusion model has comparable performance to GAN results and does not suffer from unstable training and mode collapse. Several approaches employing diffusion models demonstrate the potential of this method, such as in 2D galaxy images (Smith et al., 2022) and in 2D simulations (Schanz et al., 2023) using filter-boosted tricks. However, their work is limited to 2D cases only.
@@
In this work, we introduce an emulator framework which can adjust the SR outputs to their HR counterparts by conditioning on specific high-resolution initial conditions (HRICs). These specific HRICs have the same large-scale modes as the LR $N$-body simulations that are used to generate SR outputs. Previous studies have also utilized emulators, such as Kodi Ramanah et al. (2020b), which combines LR and HRICs for super-resolution tasks on 3D density fields. Similar to our approach, their model does not incorporate stochasticity. Furthermore, Jamieson et al. (2023) uses a neural network to map the linear approximation at $z=0$ to full $N$-body simulations under various cosmological parameters, using a Lagrangian description. To achieve accurate SR emulation of the HR simulations at small scales, we find that using adversarial training techniques is crucial for correcting small-scale discrepancies. In this study, we explore emulators trained with this technique (hereafter referred to as "Emulator"), and compare them to our previous work in Zhang et al. (2023) (hereafter referred to as "SR").
@@
The paper is structured as follows. In Section 2, we discuss the generation of data sets and summarize the details of our neural networks and the training process. In Section 3 we show the performance of our emulator, specifically visual validation in 3.1, Fourier mode cross-correlations in 3.2 and halo catalog in 3.3. In Section 4, we discuss our findings, and we summarize our results and conclude in Section 5. We have also included the results of the training without discriminator guidance (hereafter referred to as Emulator w/o DG) in the appendix A.
@@ 2.1 Dataset
The $N$-body simulation is a widely used tool for understanding the nonlinear evolution of cosmic structures. Given a fixed set of cosmological parameters (in this work, we use the WMAP9 cosmology, as prescribed by Hinshaw et al., 2013), this method involves dividing the mass distribution into $N$ equal-mass particles. We run both LR and HR simulations using MP-Gadget [^1] https://github.com/MP-Gadget/MP-Gadget, a highly parallel code that solves the gravitational force using the TreePM method (Bagla, 2002). Long-range forces are calculated on a particle mesh in Fourier space, and short-range forces are computed through a hierarchical tree structure. Our LR and HR simulation tasks follow the Lagrangian description, where particles are tracers of both displacement and velocity fields. The displacement of an individual particle is:
@@
$\mathbf{v}_{i}=\dot{\mathbf{x}_{i}}=\dot{\mathbf{d}_{i}},$ (2)
where the dots denote time derivatives. The initial conditions for our simulations are generated at $z=99$ using a power spectrum calculated by CLASS (Lesgourgues, 2011). First-order Lagrangian perturbation theory (Zel’dovich, 1970; Crocce et al., 2006) is used to initialize particle positions and velocities. The definition of initial displacement $\mathbf{d}_{init,i}$ and velocity $\mathbf{v}_{init,i}$ for particles is similar to Eq. 1 and 2. Cosmological parameters follow the WMAP9 cosmology (Hinshaw et al., 2013), with matter density $\Omega_{\rm m}=0.2814$, dark energy density $\Omega_{\Lambda}=0.7186$, baryon density $\Omega{\rm b}=0.0464$, power spectrum normalization $\sigma_{8}=0.82$, spectral index $n_{s}=0.971$, and Hubble parameter $h=0.697$. The mean spatial separation of the dark matter particles is used to determine the gravitational softening length, which is set to $1/30$ of this value. Our training dataset is comprised of 8 dark matter only simulations, each with LR and HR simulation pairs. Each LR-HR simulation pair has $64^{3}$ and $512^{3}$ particles, respectively, with box size $100h^{-1}{\rm Mpc}$. The mass resolution is $m_{\mathrm{DM}}=2.98\times 10^{11}{h^{-1}M_{\odot}}$ for LR, and $m_{\mathrm{DM}}=5.8\times 10^{8}{h^{-1}M_{\odot}}$ for HR. Each pair has a unique random seed for its initial condition, so the initial conditions differ for eight simulations. In addition, we also create a test dataset with 3 LR and HR simulation pairs. Each pair consists of 60 snapshots from $z=1$ to $z=0$, with the same cosmological parameters as the training set but different initial condition random seeds. During data pre-processing, the particle displacement and velocity fields are divided into cubical chunks. These fields are then concatenated channel-wise, as they share the same spatial resolution. The same procedure is applied to the HRICs.
@@ 2.2 Model Architecture and General Training Process
Figure 1: Upper Panel : A schematic representation of the main components of the model and the overall training process. Lower Panel : The architecture of our U-Net model and the detailed structure of the Residual Blocks. The Res up Block and Res Down Block represent the up-sampling and down-sampling components, respectively, while the Res Blocks do not include up-sampling or down-sampling operations. For a more detailed description, refer to Section 2.2.
@@
Our model employs a U-Net architecture, named for its U-shaped design, originally developed for image segmentation (Ronneberger et al., 2015) and adapted here for 3D tasks. U-Nets have also been applied to super-resolution and style transfer tasks (Johnson et al., 2016), as well as to image to image translation, image uncropping, and more (see discussion in Saharia et al., 2021, for a review). Recent diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) also use U-Nets for denoising. The U-Net architecture consists of an encoder and a decoder connected by a bottleneck part. The U-Net architecture consists of an encoder, a bottleneck, and a decoder. The encoder progressively downsamples the input towards the bottleneck, while the decoder upsamples the data back to its original resolution. "Skip-connections" link the encoder and decoder at corresponding resolutions, preserving spatial information throughout the network.
@@
The upper panel of Figure 1 illustrates the overall layout of our architecture and training process. The model’s redshift encoder consists of a multi-layer perceptron designed to extract redshift information, enabling the generation of outputs at multiple epochs. The initial condition and SR realization are concatenated channel-wise and then fed into the U-Net model. The output simulation is then compared to the target HR simulation using trainable discriminator. During training, backpropagation is performed using Pytorch, with a combined loss function employed to optimize particle-level details. The lower panel of Figure 1 details each component’s role. The encoder part extracts features from the input and downsamples them towards the bottleneck. This forces the model to learn the key information from the input and embed it into a low-dimensional space. Starting from the input resolution, the decoder downsamples the learned feature until it reaches the bottleneck blocks. Each decoder block at the same resolution consists of two residual blocks. We follow BigGAN’s (Brock et al., 2018) residual block structure since it has been proven to have good performance and can be scaled to different depths. All residual blocks contain two branches, one “skip” branch with a 3D size 1 convolutional block, the other “main” branch, which first normalizes the input with Group Normalization (Wu & He, 2018), then passes through a Sigmoid Linear Unit (SiLU) activation function (Elfwing et al., 2017). The second group normalization layer is applied to normalize the output from a 3D size 3 convolutional block. Finally, after one more activation function and one 3D size 3 convolutional block, the output from the “main” branch is added to the output from the ”skip” branch. The Up/Down layer denotes the upsampling or downsampling procedure, only applied in Res Up/Down Blocks. The decoder part decodes and upsamples the features from the bottleneck and the ’skip’ connections from the encoder. The information from the encoder is passed to the decoder at multiple resolution levels, to maintain information at different levels and avoid vanishing gradient issues. In this model, no stochasticity is injected.
@@
The goal of our model is to adjust the fields generated by the stochastic SR model to match as closely as possible the output of a specific HR simulation, conditioned on the HRICs. Due to the limitations of GPU memory, we first split the SR field and HRICs into equal size chunks, then concatenate them in a channel-wise fashion since they have the same spatial dimensions. Both chunks pass through Emulator, and are then compared to the real HR simulations generated from the same HRICs. To evaluate our emulator’s output, we use the Charbonnier loss function (Barron, 2017), which is defined as
@@
Using a discriminator to guide the training process (i.e., through adversarial loss) is a common technique in image generation and image-to-image translation tasks, as it enhances the perceptual quality of the output. We adopt the same definition of adversarial loss, $\mathcal{L}_{Adv}$ as in previous SR works (Li et al., 2021; Ni et al., 2021; Zhang et al., 2023). The total loss for our emulator network is:
@@
The parameter $\lambda$ serves as a weight to balance the contributions of the different loss terms, and its value is empirically set to $3\times 10^{-2}$. This value was determined through experimentation to achieve an optimal trade-off between accuracy and perceptual quality in the outputs.